OpenStudy (anonymous):

When air expands adiabatically (without gaining or losing heat), its pressure P and volume V are related by the equation PV^1.4=C, where C is a constant. Suppose that at a certain instant the volume is 400cm^3 and the pressure is 80kPa and is decreasing at a rate of 10kPa/min. At what rate is the volume increasing at this instant?

OpenStudy (anonymous):

pv^1.4=constant, so we will look at event 1, being the starting of time = 0 and event 2 of time = 10 min. since the equation is constant, you can think of it like this: event 1 = event 2 or pv^1.4=pv^1.4 so plug in all the knows for event 1, which is volume is 400cm^3 and the pressure is 80kPa. 80*400^1.4= pv^1.4, now plug in the info for event 2 80*400^1.4= (80-10)v^1.4 now solve for v

OpenStudy (anonymous):

v = 440, so that means in 10 minutes, the volume went from 400 to 440, so how much did it increase in 10 minutes?..